OpenStudy (anonymous):

A child tosses five quarters in the air. Each quarter lands either falls so that “heads” is visible or “tails” is visible. What is the probability that more than one quarter lands on tails?

OpenStudy (anonymous):

compute the probability that none land head, subtract the result from 1 since "at least one" means "not none"

OpenStudy (anonymous):

ok i lies, it says more than one" so 2 or 3 or 4 or 5

OpenStudy (anonymous):

compute the probability you get none or one then subtract those from 1

OpenStudy (anonymous):

\[\large 1-\left(\frac{1}{2}\right)^5-5\times \left(\frac{1}{2}\right)^5\]

OpenStudy (anonymous):

@satellite73 dont mean to bother you but are you done helping me out?

OpenStudy (anonymous):

@zepdrix hey man, are you available to give me some help?

OpenStudy (anonymous):

@jim_thompson5910 hey jim, do you know anything about "two's compliment?"

jimthompson5910 (jim_thompson5910):

what's your question?

OpenStudy (anonymous):

Ok, just a moment

OpenStudy (anonymous):

1. 0001110001110001 2. 1110001110001110 + 1 1110001110001111

OpenStudy (anonymous):

oh, wait a minute. i think i just saw what i did wrong... i wasn't supposed to flip it before converting it because it's positive already

OpenStudy (anonymous):

have another question!

OpenStudy (anonymous):

Earlier in this thread Satelite gave me an equation that solves the original post. I don't understand what the parts of it mean 1-(1/2)^5 - 5x(1/2)^5

OpenStudy (anonymous):

I'm guessing that (1/2) is the probability for either side. ^5 is because we have 5 flips. but, why the 1 and the 5? if i wanted to say "more than two flips would come up heads" which number would i change?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one moment, still thinking

jimthompson5910 (jim_thompson5910):

The probability of getting heads is 1/2 The probability of getting all five heads is (1/2)^5 the complement of this event is having at least one tail, so that would be 1 - (1/2)^5

OpenStudy (anonymous):

you mean that 5(1/2)^5 is the probability of 5 heads?

jimthompson5910 (jim_thompson5910):

I mean (1/2)^5 is the probability of getting 5 heads in a row

OpenStudy (anonymous):

okay, i can see that now. but, why does the -5 come in?

jimthompson5910 (jim_thompson5910):

not sure but I'm still thinking

jimthompson5910 (jim_thompson5910):

I think it has to do with the probability of getting exactly one tail

OpenStudy (anonymous):

this is important concept for me to grasp tonight cause this is one of the chapters on my test tomorrow :(

jimthompson5910 (jim_thompson5910):

I'm guessing you're being tested on the binomial probability theorem?

OpenStudy (anonymous):

Yes, i think so. I just looked through my book, though and i don't see that exact term.

jimthompson5910 (jim_thompson5910):

have you seen this formula before \[\Large \left(_n C _k\right)*p^k*(1-p)^{n-k}\]

OpenStudy (anonymous):

no, but i'm trying to find it

OpenStudy (anonymous):

is that what satelite used?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

what I posted is the binomial probability formula. It's used to answer questions like "what is the probability of getting exactly one tail?"

OpenStudy (anonymous):

there are 4 variables in your formula :(

OpenStudy (anonymous):

How would I plug my problem into this monster?

jimthompson5910 (jim_thompson5910):

n = 5 because there are 5 quarters (5 trials) p = 0.5 is the probability of getting tails k = 1 if we are asking "what is the probability of getting exactly one tail?"

jimthompson5910 (jim_thompson5910):

C is NOT a variable. It stands for the combination formula \[\Large _nC_r = \frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

I know that notation

jimthompson5910 (jim_thompson5910):

you can use a calculator to compute the combination or you can use Pascal's Triangle

OpenStudy (anonymous):

Wow, so what Satelite did was wrong?

OpenStudy (anonymous):

that seems like a lot of math

jimthompson5910 (jim_thompson5910):

P(more than 1 tail) = 1 - P(no tails) - P(exactly one tail) we already calculated `P(no tails)` when we computed (1/2)^5. We could have used the binomial theorem but no need here you just need `P(exactly one tail)` now

OpenStudy (anonymous):

what would r be?

OpenStudy (anonymous):

@satellite73 hey man, could you explain the formula you gave me earlier to answer this question?

jimthompson5910 (jim_thompson5910):

in the case of exactly 1 tail, r = 1